# Scaling Laws for Task-Optimized Models of the Primate Visual Ventral Stream

### v2

@@ 4.3 More Data Is Better Than More Parameters
Figure 3: Scaling Dataset Size. a) Training on larger datasets enhances brain alignment. The alignment scaling curves derived from ImageNet and EcoSet closely estimate the alignment achieved when using ImageNet21k. In contrast, datasets with specialized image distributions—such as Places365—fall below the alignment scaling laws established by these generalist datasets. b) Model families with weaker inductive bias start at a lower alignment and require more data to improve.
@@
In contrast, expanding the model size without proportionally increasing the training data results in steeper diminishing returns in alignment performance. Larger models rapidly reach a point where additional parameters do not translate into meaningful improvements. Fig. 2c estimates a saturation level of $0.44$ by scaling model sizes with all samples of training data whereas Fig. 3a predicts maximum alignment of $0.48$ and $0.50$ for ImageNet and Ecoset respectively. This indicates that scaling training datasets overall improves brain alignment better than models scaling. Furthermore, Fig. 4b demonstrates that larger models of the same architecture family require much more samples to achieve the same level of alignment.
@@ 4.4 Ordered effect of scale on alignment
Figure 5: Graded Effect of Scale across Cortical Hierarchy. a) Alignment as a function of training compute across different brain regions. Group 1 contains most models except those with low inductive bias (Group 2; ConvNeXt, ViT). b) Alignment gain per region, defined as $A{10}^{\alpha}$. Regions higher in the cortical hierarchy show greater benefits from increased compute (Behavior $>$ IT $>$ V4 $>$ V2 $>$ V1).
@@ 5 Discussion
Platonic representations.
Our results are consistent with the view that representations in deep neural network models are converging. This observation has been likened to Plato’s concept of an ideal reality (Huh et al., 2024). Our findings support this view in that neural network models at scale tend to converge toward similar brain and behavioral alignment, regardless of their initial architectural differences. However, we find that the representational convergence in current models is distinct from representations in the brain’s visual ventral stream. Intriguingly, behavioral choices on the other hand seem to converge to human-like core object recognition, suggesting that there is no unique solution to human-like behavior. Even models with differing inductive biases, such as ConvNeXt and ViT (which start with poorer neural alignment), eventually converge to the same behavioral, and, saturating brain alignment as models like ResNet and EfficientNet. Building models with representations that are consistent with the brain might thus require substantial changes to current architectures and training approaches.
@@
This disparity is further highlighted by the correlation between task performance and alignment depicted in Figure 6. Behavioral alignment closely tracks validation accuracy, improving hand-in-hand as models become more accurate. Consistent with prior work(Schrimpf et al., 2018; Linsley et al., 2023), neural alignment eventually saturates, indicating that factors other than task performance influence neural alignment.
@@
We assessed whether alternative training paradigms can overcome the limitations observed in neural alignment under supervised learning. Figure 7a illustrates the scaling of alignment as a function of compute spent during self-supervised training of ResNet models using SimCLR Chen et al. (2020) on ImageNet. The results confirm the trends observed in supervised training: behavioral alignment continues to improve with increased compute, following a strong power-law relationship, while neural alignment approaches a saturation point. This consistency suggests that the saturation in neural alignment is not exclusive to supervised learning but may be inherent to the models or datasets employed.
The region-specific breakdown (as illustrated in Figure 13 of the appendix) further reinforces this observation. Even in a self-supervised learning context, higher-level visual areas like IT and behavioral outputs demonstrate more pronounced improvements with increased compute, while early visual areas like V1 and V2 show minimal gains. This suggests that the hierarchical nature of neural alignment is a fundamental characteristic that transcends specific training methodologies.
Additionally, we explored the impact of adversarial fine-tuning on alignment performance. In Figure 7b, ResNet models trained on subsets of ImageNet were fine-tuned adversarially for 10 epochs using the Fast Gradient Sign Method (FGSM) (Goodfellow et al., 2015; Wong et al., 2020). Importantly, the scaling curves were estimated solely from the non-adversarial runs, yet the adversarially fine-tuned models exhibited improvements along these existing scaling curves. This indicates that adversarial training can enhance alignment without deviating from the established scaling behavior.
@@
Further analysis in Figure 10 of the appendix confirms that this initial high alignment is due to the strong inductive biases present in fully convolutional networks. These biases enable the models to start with representations already well-suited for neural alignment. Figure 11 in the appendix reinforces this finding by demonstrating that models with strong inductive biases achieve higher initial alignment compared to architectures like ConvNeXt and ViT, which have weaker inductive biases.
@@
Our investigation reveals that the type of learning signal plays a crucial role in the dynamics of alignment during training. Figure 7d illustrates the alignment trajectories of ResNet50 and ViT-S models trained on ImageNet using supervised learning, SimCLR, and DINO (Caron et al., 2021) methods. Notably, the ViT-S model requires significantly more training steps to achieve the same level of alignment under supervised learning compared to when trained with self-supervised objectives like DINO and SimCLR. In contrast, the ResNet50 model, which possesses strong inductive biases due to its convolutional architecture, exhibits relatively consistent alignment dynamics across different learning signals. This robustness implies that models with strong inductive biases are less affected by the choice of training objective, whereas architectures like ViT-S benefit more substantially from rich, self-supervised feedback to achieve optimal alignment.
@@
To push neural alignment beyond current saturation levels, future research should explore adversarial training methods that encourage models to learn more robust, brain-like representations. Leveraging biologically inspired architectures such as VOneNets (Dapello et al., 2020) may lead to more compute-efficient models achieving higher neural alignment without extensive scaling. Additionally, investigating co-training with brain data—integrating neural recordings directly into the training process—could enhance both neural and behavioral alignment, paving the way for more accurate and efficient brain-like models.
Figure 7: Alignment Scaling of Alternative Training Strategies. a) Unsupervised: ResNet models trained with SimCLR improve behavioral alignment with compute, while neural alignment saturates. b) Adversarial Robustness: Fine-tuning ResNets with adversarial training (FGSM) enhances alignment along the scaling curve. c) Architectural Prior: ResNets and EfficientNets exhibit higher initial alignment in early phases of training, due to strong inductive biases, unlike ConvNeXt and ViT. d) Alignment dynamics vary with training objectives, but converge to the same alignment saturation.
@@ Appendix A Implementation Details
Our experiments are conducted using the PyTorch framework (Paszke et al., 2019), with Composer (Team, 2021) employed as the GPU orchestration tool to efficiently manage computational resources.
For image augmentations, we leverage the Albumentations Buslaev et al., 2020 library due to its rich set of augmentation techniques, which are crucial for enhancing model robustness and preventing overfitting. In experiments involving self-supervised learning, we use the Lightly Susmelj et al. (2020) library to facilitate the implementation of self-supervised losses, augmentations, and model heads. This library streamlines the process of setting up models for SimCLR and DINO training methods.
To generate adversarial examples for adversarial fine-tuning, we employ the Torchattacks library Kim (2020). Specifically, we use the Fast Gradient Sign Method (FGSM) to create perturbations that challenge the models, aiming to enhance their alignment with neural representations by exposing them to adversarial inputs.
@@ B.1 ImageNet21k-P
ImageNet-21k-P is a processed subset of the full ImageNet-21k dataset (Ridnik et al., 2021), which originally contains over 14 million images organized into more than 21,000 categories following the WordNet hierarchy. The ”P” denotes a pruned version where classes with insufficient images or noisy labels are filtered out to enhance dataset quality. This results in a refined dataset that maintains the richness of the original ImageNet-21k while improving label accuracy and image relevance. The resulting dataset contains approximately 11 million training images across 10,450 classes.
@@ B.2 WebVision-P
The WebVision dataset (Li et al., 2017) is a large-scale web image dataset designed to provide a real-world, noisy alternative to ImageNet. It originally contains over 16 million images categorized into 5,000 classes. The images are collected from the internet using queries from search engines like Google and Flickr, leading to a dataset that includes label noise, varying image resolutions, and diverse visual contexts. Due to classes with very few available samples, we processed the WebVision dataset similarly to ImageNet-21k-P to remove classes with insufficient images. The resulting dataset, which we denote as WebVision-P, contains approximately 13.5 million training images across 4,189 categories.
@@
iNaturalist Van Horn et al. (2018) contains 2.7 million photographs of organisms in their natural environments, representing 10000 species. The dataset features highly specialized fine-grained categories and natural backgrounds, offering insight into how domain-specific visual features influence alignment scaling.
@@ B.4 Places365
Places365 Zhou et al. (2017) is a large-scale scene-centric dataset containing approximately 1.8 million training images across 365 scene categories. Unlike object-centric datasets such as ImageNet, Places365 focuses on the recognition of environmental scenes, including natural landscapes, urban settings, and indoor environments. Each category includes a wide variety of images to capture the diversity within scene types.
@@
The MNIST dataset (LeCun et al., 1998) is a classic benchmark in machine learning, comprising 70,000 grayscale images of handwritten digits (0-9), each sized 28×28 pixels. To expand this dataset for more extensive experimentation, we utilize the Infinite MNIST (Infimnist) tool Loosli et al., 2007, which generates additional MNIST-like samples through data augmentation techniques. We create an extended dataset by modifying the original training dataset 19 additional times, resulting in a total of 1.2 million images. This enlarged dataset allows for a more thorough evaluation of scaling effects on the alignment.
@@ Appendix C Validation on private data
Figure 8: Public benchmarks used in this study correlates highly with private benchmarks on Brain-Score
@@ Appendix D Pretrained Models
As part of our comprehensive evaluation, we benchmarked a diverse set of pretrained models sourced from both torchvision (maintainers & contributors, 2016) and the timm (Wightman, 2019) libraries. We tested a total of 94, including ViT Dosovitskiy et al. (2021), DaViT Ding et al. (2022), LeViT Graham et al. (2021), ConvNeXt Liu et al. (2022a), MobileViT Mehta & Rastegari (2022), MaxVit Tu et al. (2022), FastViT Vasu et al. (2023). Each model varies in parameter count, training sample size, dataset source, and training objective, providing a broad spectrum for analysis.
To verify the generalizability of our findings, we conducted evaluations with these pretrained models, including larger networks like CLIP (Radford et al., 2021) and DINOv2 (Oquab et al., 2023), which are pretrained on richer and more diverse datasets such as LAION Schuhmann et al. (2021; 2022). We also compared variations of these models by examining base pretrained models alongside their fine-tuned counterparts on ImageNet, aiming to investigate the impact of fine-tuning on scaling behavior.
@@
The curves in Figure 9 closely follow the scaling patterns estimated for our trained models shown in Figure 2.c, further validating that the observed saturation is consistent across different pretraining regimes and dataset scales. This reinforces our conclusion that scaling alone is insufficient to overcome the limitations in neural alignment and highlights the need for alternative approaches to improve alignment with neural representations.
Figure 9: Alignment of pretrained vision models as a function of model parameters
@@ Appendix E Training Evolution
Figure 10: Evolution of per-region alignment throughout training. Models with stronger priors—such as ResNet and EfficientNet—exhibit higher neural alignment initially. However, the gap in representational power diminishes as more generalist models are trained on data.
Figure 11: Some untrained models demonstrate non-zero neural alignment at initialization. Nevertheless, all models start with almost zero behavioral alignment, indicating that initial neural alignment arises from architectural biases rather than learned behavior.
@@ Appendix F Effect of Training Objective
Figure 12: Effect of training objective on alignment
Figure 13: Scaling of SimCLR training across regions
@@ Appendix G Region-wise scatter plots
Figure 14: Region vs. Region Comparisons: This figure shows how the alignment scores for each brain region correlate with those of other regions. The diagonal plots illustrate the relationship between the alignment score of each region and the validation accuracy on ImageNet and EcoSet datasets.

### v3

@@ 3.3 More Data Is Better Than More Parameters
Figure 3: Scaling Dataset Size. a) Training on larger datasets enhances brain alignment. The alignment scaling curves derived from ImageNet and EcoSet closely estimate the alignment achieved when using ImageNet21k. In contrast, datasets with specialized image distributions—such as Places365—fall below the alignment scaling laws established by these generalist datasets. In the extreme case of handwritten digits (infiMNIST), the impact of training is minuscule on the alignment. b) Model families with weaker inductive bias start at a lower alignment and require more data to improve.
@@
In contrast, expanding the model size without proportionally increasing the training data results in steeper diminishing returns in alignment performance. Larger models rapidly reach a point where additional parameters do not translate into meaningful improvements. Fig. 2c estimates a saturation level of $0.44$ by scaling model sizes with all samples of training data whereas Fig. 3a predicts maximum alignment of $0.46$ and $0.48$ for ImageNet and Ecoset respectively. This indicates that scaling training datasets overall improves brain alignment better than models scaling. Furthermore, Fig. 4b demonstrates that larger models of the same architecture family require much more samples to achieve the same level of alignment.
@@ 3.4 Ordered effect of scale on alignment
Figure 5: Graded Effect of Scale across Cortical Hierarchy. a) Alignment as a function of training compute across different brain regions. Group 1 contains most models except those with low inductive bias (Group 2; ConvNeXt, ViT). b) Alignment gain per region, defined as $A{10}^{\alpha}$. Regions higher in the cortical hierarchy show greater benefits from increased compute (Behavior >> IT >> V4 >> V2 >> V1).
@@ 4 Discussion
We find a saturation of neural alignment under current modeling approaches, consistent with trends reported in prior work [linsley2023worseit, conwell2024, muttenthaler2023human]. Critically, our results reveal a disconnect between neural and behavioral alignment: while behavioral alignment continues to improve with increased scale, neural alignment plateaus. By quantifying scaling laws across model families and data regimes, we show that improvements in brain alignment are more efficiently achieved by increasing dataset size rather than model parameters. These findings offer concrete guidance for developing brain-like models more effectively, emphasizing the importance of dataset diversity and biologically inspired architectural priors over brute-force model scaling.
@@
This disparity is further highlighted by the correlation between task performance and alignment depicted in Figure 6. Behavioral alignment closely tracks validation accuracy, improving hand-in-hand as models become more accurate. Consistent with prior work[SchrimpfKubilius2018BrainScore, linsley2023worseit], neural alignment eventually saturates, indicating that factors other than task performance influence neural alignment.
@@
We assessed whether alternative training paradigms can overcome the limitations observed in neural alignment under supervised learning. Figure 7a illustrates the scaling of alignment as a function of compute spent during self-supervised training of ResNet models using SimCLR [chen2020simclr] on ImageNet. The results confirm the trends observed in supervised training: behavioral alignment continues to improve with increased compute, following a strong power-law relationship, while neural alignment approaches a saturation point. This consistency suggests that the saturation in neural alignment is not exclusive to supervised learning but may be inherent to the models or datasets employed.
The region-specific breakdown (as illustrated in Supp. Fig. S6) further reinforces this observation. Even in a self-supervised learning context, higher-level visual areas like IT and behavioral outputs demonstrate more pronounced improvements with increased compute, while early visual areas like V1 and V2 show minimal gains. This suggests that the hierarchical nature of neural alignment is a fundamental characteristic that transcends specific training methods.
Additionally, we explored the impact of adversarial fine-tuning on alignment performance. In Figure 7b, ResNet models trained on subsets of ImageNet were fine-tuned adversarially for 10 epochs using the Fast Gradient Sign Method (FGSM) [Goodfellow2015fgsm, Wong2020ffgsm]. Importantly, the scaling curves were estimated solely from the non-adversarial runs, yet the adversarially fine-tuned models exhibited improvements along these existing scaling curves. This indicates that adversarial training can enhance alignment without deviating from the established scaling behavior.
Figure 7: Alignment Scaling of Alternative Training Strategies. a) SimCLR Training: ResNet models trained with SimCLR improve behavioral alignment with compute, while neural alignment saturates. b) Adversarial Robustness: Fine-tuning ResNets with adversarial training (FGSM) enhances alignment along the scaling curve. c) Supervised Training Dynamics: ResNets and EfficientNets exhibit higher initial alignment in early phases of training, due to strong inductive biases, unlike ConvNeXt and ViT. d) Self-supervised Training (variants):Alignment dynamics vary with training objectives, but converge to the same alignment saturation.
@@
Further analysis in Figure S3 confirms that this initial high alignment is due to the strong inductive biases present in fully convolutional networks. These biases enable the models to start with representations already well-suited for neural alignment. Figure S4 reinforces this finding by demonstrating that models with strong inductive biases achieve higher initial alignment compared to architectures like ConvNeXt and ViT, which have weaker inductive biases.
@@
Our investigation reveals that the type of learning signal plays a crucial role in the dynamics of alignment during training. Figure 7d illustrates the alignment trajectories of ResNet50 and ViT-S models trained on ImageNet using supervised learning, SimCLR, and DINO [caron2021dinov1] methods. Notably, the ViT-S model requires significantly more training steps to achieve the same level of alignment under supervised learning compared to when trained with self-supervised objectives like DINO and SimCLR. In contrast, the ResNet50 model, which possesses strong inductive biases due to its convolutional architecture, exhibits relatively consistent alignment dynamics across different learning signals. This robustness implies that models with strong inductive biases are less affected by the choice of training objective, whereas architectures like ViT-S benefit more substantially from rich, self-supervised feedback to achieve optimal alignment.
@@
While transformer-based architectures like ViTs dominate in many ML tasks due to their flexibility and strong task performance, our findings highlight a clear dissociation between behavioral and neural alignment. Convolutional models exhibit significantly better neural alignment, especially in early training and low-data regimes—suggesting that biologically inspired inductive biases play a unique role in approximating cortical representations. Interestingly, we also show that recurrence (e.g., CORnet) offers benefits in sample efficiency, but its advantage diminishes with more data. These observations motivate hybrid architectures that blend convolutional, recurrent, and transformer components.
@@
To push neural alignment beyond current saturation levels, future research should explore adversarial training methods that encourage models to learn more robust, brain-like representations. Leveraging biologically inspired architectures such as VOneNets [vonenet] may lead to more compute-efficient models achieving higher neural alignment without extensive scaling. Additionally, investigating co-training with brain data—integrating neural recordings directly into the training process—could enhance both neural and behavioral alignment, paving the way for more accurate and efficient brain-like models.
Acknowledgements
This work was partly supported by the Swiss National Science Foundation (SNSF) through the Spark project CRSK-3_228579. We thank the members of the EPFL NeuroAI Lab for their valuable discussions and feedback throughout the project.
Software and Data
We open-source our training and analysis code, along with benchmark results and model checkpoints from our model zoo. All resources are available at: https://github.com/epflneuroailab/scaling-primate-vvs.
Impact Statement
This work seeks to advance our understanding of how scaling model size and dataset size influences alignment with primate visual processing. While the primary goal is scientific—improving models of the visual cortex—it also intersects with broader concerns about resources and impact. Training large-scale models has nontrivial computational costs and environmental implications. At the same time, more accurate brain-like models could yield downstream benefits for neuroscience and machine learning applications (for instance, improved medical imaging or more robust computer vision systems). We do not foresee direct ethical risks specific to this research beyond the usual resource considerations in large-scale model training, and we have used existing public neurophysiological and behavioral datasets.
@@ Appendix A Implementation Details
Our experiments are conducted using the PyTorch framework [pytorch], with Composer [mosaicml2022composer] employed as the GPU orchestration tool to efficiently manage computational resources.
For image augmentations, we leverage the Albumentations Albumentations library due to its rich set of augmentation techniques, which are crucial for enhancing model robustness and preventing overfitting. In experiments involving self-supervised learning, we use the Lightly [susmelj2020lightly] library to facilitate the implementation of self-supervised losses, augmentations, and model heads. This library streamlines the process of setting up models for SimCLR and DINO training methods.
To generate adversarial examples for adversarial fine-tuning, we employ the Torchattacks library [kim2020torchattacks]. Specifically, we use the Fast Gradient Sign Method (FGSM) to create perturbations that challenge the models, aiming to enhance their alignment with neural representations by exposing them to adversarial inputs.
@@ B.1 ImageNet21k-P
ImageNet-21k-P is a processed subset of the full ImageNet-21k dataset [ridnik2021imagenet21k], which originally contains over 14 million images organized into more than 21,000 categories following the WordNet hierarchy. The ”P” denotes a pruned version where classes with insufficient images or noisy labels are filtered out to enhance dataset quality. This results in a refined dataset that maintains the richness of the original ImageNet-21k while improving label accuracy and image relevance. The resulting dataset contains approximately 11 million training images across 10,450 classes.
@@ B.2 WebVision-P
The WebVision dataset [webvision] is a large-scale web image dataset designed to provide a real-world, noisy alternative to ImageNet. It originally contains over 16 million images categorized into 5,000 classes. The images are collected from the internet using queries from search engines like Google and Flickr, leading to a dataset that includes label noise, varying image resolutions, and diverse visual contexts. Due to classes with very few available samples, we processed the WebVision dataset similarly to ImageNet-21k-P to remove classes with insufficient images. The resulting dataset, which we denote as WebVision-P, contains approximately 13.5 million training images across 4,189 categories.
@@
iNaturalist [horn2018inaturalist] contains 2.7 million photographs of organisms in their natural environments, representing 10000 species. The dataset features highly specialized fine-grained categories and natural backgrounds, offering insight into how domain-specific visual features influence alignment scaling.
@@ B.4 Places365
Places365 [zhou2017places365] is a large-scale scene-centric dataset containing approximately 1.8 million training images across 365 scene categories. Unlike object-centric datasets such as ImageNet, Places365 focuses on the recognition of environmental scenes, including natural landscapes, urban settings, and indoor environments. Each category includes a wide variety of images to capture the diversity within scene types.
@@
The MNIST dataset [mnist] is a classic benchmark in machine learning, comprising 70,000 grayscale images of handwritten digits (0-9), each sized 28×28 pixels. To expand this dataset for more extensive experimentation, we utilize the Infinite MNIST (Infimnist) tool infimnist, which generates additional MNIST-like samples through data augmentation techniques. We create an extended dataset by modifying the original training dataset 19 additional times, resulting in a total of 1.2 million images. This enlarged dataset allows for a more thorough evaluation of scaling effects on the alignment.
B.6 CIFAR10
CIFAR-10 [cifar10] is a widely used benchmark of 60,000 low-resolution (32×32) color images divided evenly into 10 object classes. The dataset comprises 50,000 training images and 10,000 test images, with 6,000 samples per class. To match our scaling protocol, we created class-balanced subsets by sampling $d\,\in\,\{10,\,30,\,100,\,300,\,1000,\,3000\}$ images per class. Because CIFAR-10 is two orders of magnitude smaller than our other baselines, in addition to our standard 100-epoch runs we also trained models on the full dataset for extended durations of 250, 1000, and 2500 epochs to assess convergence and scaling effects.
@@ Appendix C Validation on private data
Figure S1: Public benchmarks used in this study correlates highly with private benchmarks on Brain-Score
@@ Appendix D Pretrained Models
As part of our comprehensive evaluation, we benchmarked a diverse set of pretrained models sourced from both torchvision [torchvision2016] and the timm [rw2019timm] libraries. We tested a total of 94, including ViT [vit2021], DaViT [davit2022], LeViT [levit2021], ConvNeXt [convnetx2022], MobileViT [mobilevit2022], MaxVit [maxvit2022], FastViT [fastvit2023]. Each model varies in parameter count, training sample size, dataset source, and training objective, providing a broad spectrum for analysis.
To verify the generalizability of our findings, we conducted evaluations with these pretrained models, including larger networks like CLIP [radford2021clipopenai] and DINOv2 [oquab2023dinov2], which are pretrained on richer and more diverse datasets such as LAION [laion400m2021, laion5b2022]. We also compared variations of these models by examining base pretrained models alongside their fine-tuned counterparts on ImageNet, aiming to investigate the impact of fine-tuning on scaling behavior.
@@
The curves in Figure S2 closely follow the scaling patterns estimated for our trained models shown in Figure 2.c, further validating that the observed saturation is consistent across different pretraining regimes and dataset scales. This reinforces our conclusion that scaling alone is insufficient to overcome the limitations in neural alignment and highlights the need for alternative approaches to improve alignment with neural representations.
Figure S2: Alignment of pretrained vision models as a function of model parameters
@@ Appendix E Training Evolution
Figure S3: Evolution of per-region alignment throughout training. Models with stronger priors—such as ResNet and EfficientNet—exhibit higher neural alignment initially. However, the gap in representational power diminishes as more generalist models are trained on data.
Figure S4: Some untrained models demonstrate non-zero neural alignment at initialization. Nevertheless, all models start with almost zero behavioral alignment, indicating that initial neural alignment arises from architectural biases rather than learned behavior.
@@ Appendix F Effect of Training Objective
Figure S5: Effect of training objective on alignment
Figure S6: Scaling of SimCLR training across regions
@@ Appendix G Region-wise scatter plots
Figure S7: Region vs. Region Comparisons: This figure shows how the alignment scores for each brain region correlate with those of other regions. The diagonal plots illustrate the relationship between the alignment score of each region and the validation accuracy on ImageNet and EcoSet datasets.